# Impact of Dataset Properties on Membership Inference Vulnerability of Deep Transfer Learning

### v3

@@ C.2 Proof of Lemma 2
We abuse notations by denoting probabilities and expectations over sampling $\mathcal{D}_{\mathrm{target}}\sim{\mathbb{D}}^{|\mathcal{D}|}$ and $(z,y_{z})\sim{\mathbb{D}}$ by subscripts $\mathcal{D}_{\mathrm{target}}$ and $z$. Recall that
$\textsc{tpr}_{\mathrm{RMIA}}(x)$ $=\Pr_{\mathcal{D}_{\mathrm{target}}}\left(\Pr_{z}\left(\mathrm{LR}(x,z)\geq\gamma\right)\geq\beta\mid(x,y_{x})\in\mathcal{D}_{\mathrm{target}}\wedge(x,y_{x})\notin\mathcal{D}_{\mathrm{target}}\right)$ (A43)
$\textsc{fpr}_{\mathrm{RMIA}}(x)$ $=\Pr_{\mathcal{D}_{\mathrm{target}}}\left(\Pr_{z}\left(\mathrm{LR}(x,z)\geq\gamma\right)\geq\beta\mid(x,y_{x})\notin\mathcal{D}_{\mathrm{target}}\wedge(x,y_{x})\in\mathcal{D}_{\mathrm{target}}\right),$ (A44)
where
$\mathrm{LR}(x,z)=\frac{\mathcal{N}(t_{x};\hat{\mu}_{x,x},\hat{\sigma}_{x}^{2})\mathcal{N}(t_{z};\hat{\mu}_{x,z},\hat{\sigma}_{z}^{2})}{\mathcal{N}(t_{x};\hat{\mu}_{z,x},\hat{\sigma}_{x}^{2})\mathcal{N}(t_{z};\hat{\mu}_{z,z},\hat{\sigma}_{z}^{2})}.$ (A45)
We have
$\lambda:=$ $\log(\mathrm{LR}(x,z))$ (A46)
$=$ $\log\left(\frac{\frac{1}{\sqrt{2\pi}\hat{\sigma}_{x}}\exp\left(-\frac{1}{2}\left(\frac{t_{x}-\hat{\mu}_{x,x}}{\hat{\sigma}_{x}}\right)^{2}\right)\frac{1}{\sqrt{2\pi}\hat{\sigma}_{z}}\exp\left(-\frac{1}{2}\left(\frac{t_{z}-\hat{\mu}_{x,z}}{\hat{\sigma}_{z}}\right)^{2}\right)}{\frac{1}{\sqrt{2\pi}\hat{\sigma}_{x}}\exp\left(-\frac{1}{2}\left(\frac{t_{x}-\hat{\mu}_{z,x}}{\hat{\sigma}_{x}}\right)^{2}\right)\frac{1}{\sqrt{2\pi}\hat{\sigma}_{z}}\exp\left(-\frac{1}{2}\left(\frac{t_{z}-\hat{\mu}_{z,z}}{\hat{\sigma}_{z}}\right)^{2}\right)}\right)$ (A47)
$=$ $-\frac{1}{2}\left(\frac{t_{x}-\hat{\mu}_{x,x}}{\hat{\sigma}_{x}}\right)^{2}+\frac{1}{2}\left(\frac{t_{x}-\hat{\mu}_{z,x}}{\hat{\sigma}_{x}}\right)^{2}-\frac{1}{2}\left(\frac{t_{z}-\hat{\mu}_{x,z}}{\hat{\sigma}_{z}}\right)^{2}+\frac{1}{2}\left(\frac{t_{z}-\hat{\mu}_{z,z}}{\hat{\sigma}_{z}}\right)^{2}$ (A48)
$=$ $\frac{1}{2\hat{\sigma}_{x}^{2}}(2t_{x}\hat{\mu}_{x,x}-\hat{\mu}_{x,x}^{2}-2t_{x}\hat{\mu}_{z,x}+\hat{\mu}_{z,x}^{2})+\frac{1}{2\hat{\sigma}_{z}^{2}}(2t_{z}\hat{\mu}_{x,z}-\hat{\mu}_{x,z}^{2}-2t_{z}\hat{\mu}_{z,z}+\hat{\mu}_{z,z}^{2})$ (A49)
$=$ $\frac{\hat{\mu}_{x,x}-\hat{\mu}_{z,x}}{2\hat{\sigma}_{x}^{2}}(2t_{x}-\hat{\mu}_{x,x}-\hat{\mu}_{z,x})+\frac{\hat{\mu}_{x,z}-\hat{\mu}_{z,z}}{2\hat{\sigma}_{z}^{2}}(2t_{z}-\hat{\mu}_{x,z}-\hat{\mu}_{z,z}).$ (A50)
@@
When $(x,y_{x})\in\mathcal{D}_{\mathrm{target}}$, $t_{x}=\mu_{x,x}+\sigma_{x}t$ and $t_{z}=\mu_{x,z}+\sigma_{z}t$, and thus $\lambda$ becomes
$\lambda_{x}:=$ $\frac{\hat{\mu}_{x,x}-\hat{\mu}_{z,x}}{2\hat{\sigma}_{x}^{2}}(2\mu_{x,x}+2\sigma_{x}t-\hat{\mu}_{x,x}-\hat{\mu}_{z,x})$ (A51)
$+\frac{\hat{\mu}_{x,z}-\hat{\mu}_{z,z}}{2\hat{\sigma}_{z}^{2}}(2\mu_{x,z}+2\sigma_{z}t-\hat{\mu}_{x,z}-\hat{\mu}_{z,z}).$ (A52)
Similarly, when $(x,y_{x})\notin\mathcal{D}_{\mathrm{target}}$, we have $t_{x}=\mu_{z,x}+\sigma_{x}t$ and $t_{z}=\mu_{z,z}+\sigma_{z}t$. Then $\lambda$ is now
$\lambda_{z}:=$ $\frac{\hat{\mu}_{x,x}-\hat{\mu}_{z,x}}{2\hat{\sigma}_{x}^{2}}(2\mu_{z,x}+2\sigma_{x}t-\hat{\mu}_{x,x}-\hat{\mu}_{z,x})$ (A53)
$+\frac{\hat{\mu}_{x,z}-\hat{\mu}_{z,z}}{2\hat{\sigma}_{z}^{2}}(2\mu_{z,z}+2\sigma_{z}t-\hat{\mu}_{x,z}-\hat{\mu}_{z,z})$ (A54)
$=$ $\left(\underbrace{\frac{\sigma_{x}}{\hat{\sigma}_{x}^{2}}(\hat{\mu}_{x,x}-\hat{\mu}_{z,x})+\frac{\sigma_{z}}{\hat{\sigma}_{z}^{2}}(\hat{\mu}_{x,z}-\hat{\mu}_{z,z})}_{A}\right)t$ (A55)
$+\underbrace{\frac{\hat{\mu}_{x,x}-\hat{\mu}_{z,x}}{2\hat{\sigma}_{x}^{2}}(2\mu_{z,x}-\hat{\mu}_{x,x}-\hat{\mu}_{z,x})+\frac{\hat{\mu}_{x,z}-\hat{\mu}_{z,z}}{2\hat{\sigma}_{z}^{2}}(2\mu_{z,z}-\hat{\mu}_{x,z}-\hat{\mu}_{z,z})}_{B}$ (A56)
$=$ $At+B.$ (A57)
Notice that $A$ and $B$ are functions of $z$ and independent of $t$. Also note that taking probability over $t$ corresponds to calculating probability over $\mathcal{D}_{\mathrm{target}}\sim{\mathbb{D}}^{|\mathcal{D}|}$. Thus, Since we can set $\gamma>1$, using Markov’s inequality and the triangle inequality, we have in the limit of infinitely many shadow models
$\textsc{fpr}_{\mathrm{RMIA}}(x)=$ $\Pr_{t}\left(\Pr_{z}(\lambda_{z}\geq\log(\gamma))\geq\beta\right)$ (A58)
$\leq\Pr_{t}\left(\Pr_{z}(|\lambda_{z}|\geq\log(\gamma))\geq\beta\right)$ (A59)
$\leq\Pr_{t}\left(\frac{\mathbb{E}_{z}[|\lambda_{z}|]}{\log(\gamma)}\geq\beta\right)$ (A60)
$=\Pr_{t}\left(\mathbb{E}_{z}[|\lambda_{z}|]\geq\beta\log(\gamma)\right)$ (A61)
$\leq\Pr_{t}\left(\mathbb{E}_{z}[|A|]|t|+\mathbb{E}_{z}[|B|]\geq\beta\log(\gamma)\right)$ (A62)
$=\Pr_{t}\left(|t|\geq\frac{\beta\log(\gamma)-\mathbb{E}_{z}[|B|]}{\mathbb{E}_{z}[|A|]}\right)$ (A63)
$\alpha=1-F_{|t|}\left(\frac{\beta\log(\gamma)-\mathbb{E}_{z}[|B|]}{\mathbb{E}_{z}[|A|]}\right).$ (A64)
That is,
$\frac{\beta\log(\gamma)-\mathbb{E}_{z}[|B|]}{\mathbb{E}_{z}[|A|]}=F_{|t|}^{-1}(1-\alpha).$ (A65)
Now let
$q=\lambda_{x}-\lambda_{z}=\frac{(\mu_{x,x}-\mu_{z,x})(\hat{\mu}_{x,x}-\hat{\mu}_{z,x})}{\hat{\sigma}_{x}^{2}}+\frac{(\mu_{x,z}-\mu_{z,z})(\hat{\mu}_{x,z}-\hat{\mu}_{z,z})}{\hat{\sigma}_{z}^{2}}.$ (A66)
Note that $q$ is also independent of $t$, thereby $\mathbb{E}_{z}[|q|]$ being a constant. By the similar argument using Markov’s inequality and the triangle inequality, we have
$\textsc{tpr}_{\mathrm{RMIA}}(x)=$ $\Pr_{t}\left(\Pr_{z}(\lambda_{x}\geq\log(\gamma))\geq\beta\right)$ (A67)
$\leq\Pr_{t}\left(\Pr_{z}(|\lambda_{x}|\geq\log(\gamma))\geq\beta\right)$ (A68)
$\leq\Pr_{t}\left(\frac{\mathbb{E}_{z}[|\lambda_{x}|]}{\log(\gamma)}\geq\beta\right)$ (A69)
$=\Pr_{t}\left(\mathbb{E}_{z}[|\lambda_{x}|]\geq\beta\log(\gamma)\right)$ (A70)
$\leq\Pr_{t}\left(\mathbb{E}_{z}[|A|]|t|+\mathbb{E}_{z}[|B|]+\mathbb{E}_{z}[|q|]\geq\beta\log(\gamma)\right)$ (A71)
$=\Pr_{t}\left(|t|\geq\frac{\beta\log(\gamma)-\mathbb{E}_{z}[|B|]}{\mathbb{E}_{z}[|A|]}-\frac{\mathbb{E}_{z}[|q|]}{\mathbb{E}_{z}[|A|]}\right)$ (A72)
Hence we obtain
$\textsc{tpr}_{\mathrm{RMIA}}(x)\leq 1-F_{|t|}^{-1}\left(F_{|t|}^{-1}(1-\alpha)-\frac{\mathbb{E}_{z}[|q|]}{\mathbb{E}_{z}[|A|]}\right).$ (A73)
C.3 Proof of Theorem 3
@@
Let $\mathcal{D}_{\mathrm{target}}=\{(x_{j,1},j),...,(x_{j,S},j)\}_{j=1}^{C}$. Then the LiRA score of the target $(x,y_{x})$ is
$s_{y_{x}}^{\mathrm{(in)}}$ $=\langle x,\frac{1}{S}\left(\sum_{i=1}^{S-1}x_{y_{x},i}+x\right)\rangle=\langle x,\frac{1}{S}\sum_{i=1}^{S}x_{y_{x},i}\rangle+\langle x,\frac{1}{S}(x-x_{y_{x},S})\rangle$ (A74)
$s_{y_{x}}^{\mathrm{(out)}}$ $=\langle x,\frac{1}{S}\sum_{i=1}^{S}x_{y_{x},i}\rangle,$ (A75)
respectively, when $(x,y_{x})\in\mathcal{D}_{\mathrm{target}}$ and when $(x,y_{x})\notin\mathcal{D}_{\mathrm{target}}$. Thus, in the limit of infinitely many shadow models we obtain
$\mu_{\mathrm{in}}-\mu_{\mathrm{out}}$ $=\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}}=\mathbb{E}[s_{y_{x}}^{\mathrm{(in)}}-s_{y_{x}}^{\mathrm{(out)}}]=\frac{1}{S}\langle x,x-m_{x}\rangle$ (A76)
$\sigma^{2}$ $=\operatorname{Var}(s_{y_{x}}^{\mathrm{(out)}})=\frac{1}{S}\operatorname{Var}(\langle x,x_{y_{x},i}\rangle)=\frac{1}{S}||x||^{2}s^{2}$ (A77)
$\textsc{tpr}_{\mathrm{LiRA}}(x)$ =$=\begin{cases}1-\Phi\left(\Phi^{-1}(1-\textsc{fpr}_{\mathrm{LiRA}}(x))-\frac{\mu_{\mathrm{in}}-\mu_{\mathrm{out}}}{\sigma}\right)&\text{ if }\hat{\mu}_{\mathrm{in}}>\hat{\mu}_{\mathrm{out}}\\ \Phi\left(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))-\frac{\mu_{\mathrm{in}}-\mu_{\mathrm{out}}}{\sigma}\right)&\text{ if }\hat{\mu}_{\mathrm{in}}<\hat{\mu}_{\mathrm{out}},\end{cases}$ (A78)
=$=\begin{cases}1-\Phi\left(\Phi^{-1}(1-\textsc{fpr}_{\mathrm{LiRA}}(x))-\frac{\langle x,x-m_{x}\rangle}{\sqrt{S}s||x||}\right)&\text{ if }\langle x,x-m_{x}\rangle>0\\ \Phi\left(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))-\frac{\langle x,x-m_{x}\rangle}{\sqrt{S}s||x||}\right)&\text{ if }\langle x,x-m_{x}\rangle<0,\end{cases}$ (A79)
By the symmetry of a normal distribution, this simplifies to
$\textsc{tpr}_{\mathrm{LiRA}}(x)=\Phi\left(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))+\frac{|\langle x,x-m_{x}\rangle|}{\sqrt{S}s||x||}\right),$ (A80)
where $\langle x,x-m_{x}\rangle\not=0$ by the assumption that $\hat{\mu}_{\mathrm{in}}\not=\hat{\mu}_{\mathrm{out}}$. This completes the first half of the theorem.
Now we have
$\textsc{tpr}_{\mathrm{LiRA}}(x)$ $=\Pr\left(\eta\leq\gamma_{0}+\frac{|\langle x,x-m_{x}\rangle|}{\sqrt{S}s||x||}\right),$ (A81)
$\textsc{fpr}_{\mathrm{LiRA}}(x)$ $=\Pr(\eta\leq\gamma_{0}),$ (A82)
where $\gamma_{0}$ is a tunable constant and $\eta\sim\mathcal{N}(0,1)$. Thus for large enough $S$ we have
$\textsc{tpr}_{\mathrm{LiRA}}(x)-\textsc{fpr}_{\mathrm{LiRA}}(x)$ $\approx p_{\eta}(\gamma_{0})\frac{|\langle x,x-m_{x}\rangle|}{\sqrt{S}||x||s}$ (A83)
$=\frac{1}{\sqrt{2\pi}}e^{-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))^{2}}\frac{|\langle x,x-m_{x}\rangle|}{\sqrt{S}||x||s},$ (A84)

### v4

@@ A.5 Proof of Theorem 2
Let $\mathcal{D}_{\mathrm{target}}=\{(\bm{x}_{j,1},j),...,(\bm{x}_{j,S},j)\}_{j=1}^{C}$. Then the LiRA score of the target $(\bm{x},y)$ is
$s_{y}^{\mathrm{(in)}}$ $=\langle\bm{x},\frac{1}{S}\left(\sum_{i=1}^{S-1}\bm{x}_{y,i}+\bm{x}\right)\rangle=\langle\bm{x},\frac{1}{S}\sum_{i=1}^{S}\bm{x}_{y,i}\rangle+\langle\bm{x},\frac{1}{S}(\bm{x}-\bm{x}_{y,S})\rangle$ (A36)
$s_{y}^{\mathrm{(out)}}$ $=\langle\bm{x},\frac{1}{S}\sum_{i=1}^{S}\bm{x}_{y,i}\rangle,$ (A37)
@@
respectively, when $(\bm{x},y)\in\mathcal{D}_{\mathrm{target}}$ and when $(\bm{x},y)\notin\mathcal{D}_{\mathrm{target}}$. Thus we obtain
$\mu_{\mathrm{in}}-\mu_{\mathrm{out}}$ $=\mathbb{E}[s_{y}^{\mathrm{(in)}}-s_{y}^{\mathrm{(out)}}]=\frac{1}{S}\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle$ (A38)
$\sigma^{2}$ $=\operatorname{Var}(s_{y}^{\mathrm{(out)}})=\frac{1}{S}\operatorname{Var}(\langle\bm{x},\bm{x}_{y,i}\rangle)=\frac{1}{S}\bm{x}^{T}\Sigma\bm{x}.$ (A39)
$\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})$ $=\Phi\left(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))+\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{S}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}\right),$ (A40)
where $\Phi$ is the cdf of the standard normal distribution. This completes the first half of the theorem.
Now let $\phi(u)$ denote the pdf of the standard normal distribution, and let
$r=\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{S}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}.$ (A41)
Using Taylor expansion of $\Phi(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))+r)$ around $r=0$, we have
$\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})$ $=\sum_{k=0}^{\infty}\frac{\Phi^{(k)}(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))}{k!}r^{k}$ (A42)
$=\textsc{fpr}_{\mathrm{LiRA}}(\bm{x})+\sum_{k=1}^{\infty}\frac{\Phi^{(k)}(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))}{k!}r^{k}$ (A43)
$=\textsc{fpr}_{\mathrm{LiRA}}(\bm{x})+\sum_{k=1}^{\infty}\frac{\phi^{(k-1)}(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))}{k!}r^{k}$ (A44)
$=\textsc{fpr}_{\mathrm{LiRA}}(\bm{x})+\sum_{k=1}^{\infty}\frac{(-1)^{k-1}\mathrm{He}_{k-1}(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x})))\phi(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))}{k!}r^{k}$ (A45)
$=\textsc{fpr}_{\mathrm{LiRA}}(\bm{x})+r\phi(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))\sum_{k=0}^{\infty}\frac{(-1)^{k}\mathrm{He}_{k}(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x})))}{(k+1)!}r^{k}.$ (A46)
It follows that
$\log(\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})-\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))$ (A47)
$=$ $\log r+\log\phi(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))+\log\left(\sum_{k=0}^{\infty}\frac{(-1)^{k}\mathrm{He}_{k}(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x})))}{(k+1)!}r^{k}\right)$ (A48)
$=$ $-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))^{2}+\log\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}$ (A49)
$+\log\left(\sum_{k=0}^{\infty}\frac{(-1)^{k}\mathrm{He}_{k}(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x})))}{(k+1)!}\left(\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{S}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}\right)^{k}\right)$ (A50)
$=$ $-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))^{2}+\log\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}$ (A51)
$+\log\left(1+\sum_{k=1}^{\infty}\frac{(-1)^{k}\mathrm{He}_{k}(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x})))}{(k+1)!}\left(\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{S}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}\right)^{k}\right)$ (A52)
$=$ $-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))^{2}+\log\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}+\log(1+\xi(S)),$ (A53)
where we have $\xi(S)=O(1/\sqrt{S})$. For large enough $S$, ignoring the residual term, we approximate
$\log(\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})-\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))\approx-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))^{2}+\log\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}.$ (A54)
A.6 Proof of Corollary 3
@@
$\overline{\textsc{tpr}}_{\mathrm{LiRA}}-\overline{\textsc{fpr}}_{\mathrm{LiRA}}$ $=\int_{\mathscr{D}}p(\bm{x})(\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})-\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))\mathrm{d}\bm{x}$ (A55)